OpenStudy (anonymous):

Determine the area bounded by the curve=cos x, the x-axis, and the vertical x=π and x=3π/2.

OpenStudy (amistre64):

a bounded area is usually how they start teaching you what an integral can do. \[\int^{3pi/2}_{pi}cos(x)\ dx\]

OpenStudy (underhill):

The integral of cos(x) is the antiderivative of cos(x), which is -sin(x). The solution of amistre's equation would then be \[-\sin (3\pi/2) + \sin(\pi)\] which equals 1 + 0 = 1 Therefore, the area under the Sandy's curve cos(x) bounded by the given lines is 1.